# Uncertainty Estimation by Human Perception versus Neural Models

### v1

@@ Abstract
Modern neural networks (NNs) often achieve high predictive accuracy but remain poorly calibrated, producing overconfident predictions even when wrong. This miscalibration poses serious challenges in applications where reliable uncertainty estimates are critical. In this work, we investigate how human perceptual uncertainty compares to uncertainty estimated by NNs. Using three vision benchmarks annotated with both human disagreement and crowdsourced confidence, we assess the correlation between model-predicted uncertainty and human-perceived uncertainty. Our results show that current methods only weakly align with human intuition, with correlations varying significantly across tasks and uncertainty metrics. Notably, we find that incorporating human-derived soft labels into the training process can improve calibration without compromising accuracy. These findings reveal a persistent gap between model and human uncertainty and highlight the potential of leveraging human insights to guide the development of more trustworthy AI systems.
@@ 1 Introduction
Further, real-world applications, such as medical diagnostics, autonomous driving, and fraud detection, increasingly leverage model uncertainty to enhance trustworthiness by incorporating human oversight. In these systems, models can flag predictions with high uncertainty as requiring human review. This approach allows systems to balance automation and human judgment, ensuring critical decisions are not solely reliant on potentially unreliable predictions. By raising alarms for ambiguous cases, these systems create a feedback loop where humans can validate, correct, or refine predictions, thus improving the model’s performance and fostering user trust. However, the deployment of uncertainty-based alarm systems creates a trade-off between the frequency of alarms raised and the associated costs. These costs can be reduced by aligning the model uncertainty estimation with human perception.
@@
To investigate this, we study three visual classification tasks that include human annotations of perceptual uncertainty, either through soft labels (reflecting inter-annotator disagreement) or explicit confidence scores. We compare these human-derived signals with uncertainty estimates from multiple NNs using metrics such as predictive entropy. We further test whether injecting human soft labels into training can improve the model’s alignment with human intuition and its calibration.
@@ 2 Related Work
A significant body of work has explored methods to enhance calibration of NNs. Early methods leveraged Bayesian Neural Networks (BNNs) [25, 20, 3, 37], which place a prior distribution over the model weights and infer a posterior given the data[32]. While theoretically sound, BNNs are computationally intensive and scale poorly to large models and datasets. To mitigate this, several approximate Bayesian methods have been proposed such as Monte Carlo (MC) Dropout [4] and Variational Inference (VI) [12]. Among these, MC Dropout has become the most widely used technique for uncertainty estimation in NNs [21, 1, 4]. It estimates uncertainty by applying multiple dropout masks and performing multiple stochastic forward passes, effectively simulating an ensemble of subnetworks. The resulting predictions are aggregated using statistical measures, such as predictive entropy (PE) or variance, to quantify the model’s uncertainty.
In addition to uncertainty quantification, calibration has become a key property of NNs, as they are often overconfident, even when their predictions are incorrect [8]. To address this, a range of post-hoc calibration methods have been developed, including Platt Scaling [28], Isotonic Regression [39], Temperature Scaling [8], and Beta Calibration [15], which adjust model confidence scores after training. These techniques are computationally efficient and easy to apply, but do not alter the underlying uncertainty estimates. Consequently, their effectiveness can vary significantly across architectures and datasets [19], and they tend to degrade under distributional shifts, reducing reliability in out-of-distribution (OOD) scenarios [26, 8].
@@ 3 Human and Model Uncertainty Evaluation
Therefore, this work conducts a correlational study to compare human perception of uncertainty against uncertainty estimation outputted by NNs. The objectives are twofold: i) to evaluate how well model uncertainty aligns with human perceptual uncertainty, especially in ambiguous or uncertain scenarios, and ii) to explore whether model uncertainty can be enhanced or complemented by human decision-making, identifying scenarios where human and model uncertainty assessments converge or diverge, where the model poorly estimates its uncertainty and specialize the training.
This study hypothesizes that human perceptual uncertainty strongly correlates with the uncertainty estimates generated by NNs. Specifically, we posit that as human uncertainty increases, model uncertainty will also increase. Furthermore, we hypothesize that the strength of this correlation might be dependent on task complexity, with correlations weakening as task complexity increases. Additionally, we propose that incorporating human insights can enhance the quality of model uncertainty evaluations. The findings from this research will contribute to a broader understanding of model uncertainty estimation and open pathways toward more trustworthy AI systems.
@@ 3.1 Experimental Setup, Benchmarks, and Baselines
Datasets and Models. Due to the limited resources, this work exploits publicly available datasets (namely, Cifar10-H [27], CifarN [38], and ImageNet-16H [35]) in the image recognition domain (normally used to train NNs) that include human annotations from multiple reviewers. The datasets containing human annotations have been labeled by multiple reviewers, with the number of reviews per image varying based on the benchmark used  [^1] ImageNet-16H and Cifar-N dataset received approximately four to six judgments per image, while Cifar-10H dataset has fifty annotations per image.. Further, ImageNet-16H extends the ImageNet [31] dataset by reducing the sample size to 4800 and limiting the number of classes to 16, while introducing noise in the images.
@@
To train the models, independently of the considered solution, we use stochastic gradient descent to minimize the Cross Entropy (CE) loss function using a learning rate of 0.1, a momentum of 0.9, and a batch size of 64 for all the models. To guarantee reproducibility, ensure a fair comparison, and mitigate the randomness inherent in training NNs, each method was trained using ten different seeds [17]. Throughout this study, when a subset of the dataset is required, stratified sampling techniques are applied to ensure that all classes are adequately represented and that the sample remains representative of the original dataset. Further, the models are evaluated on the same test set used for human annotations to ensure a consistent comparison between human and model uncertainty estimates. All the models and training procedures were implemented in Python3 via the Pytorch framework and trained using a single Nvidia RTX A4000.
Baselines. Furthermore, we resort to several state-of-the-art methods [16, 19, 21, 17, 8] that have been proposed to estimate model uncertainty. Moreover, one important foundation of all these works lies in the computation of uncertainty given a prediction. One of the most widely used techniques is MC Dropout [7], which provides a Bayesian approximation for uncertainty by sampling multiple dropout masks and aggregating the predictions using predictive entropy (PE). Like the human uncertainty estimation, MC Dropout combined with PE does not separate epistemic from aleatoric uncertainty, so both human and model uncertainty will reflect the total uncertainty of a prediction. In this study, we compare different baselines to compute the model uncertainty. More in detail, we evaluate both post-processing calibration methods (resorting to Isotonic Regression [39], or DEUP [16]) and uncertainty-aware training algorithms (namely, CE [40], EUAT [21], CALS [19], or deep ensemble [17]).
@@
Evaluation Metrics. We resort to the PE of the outputted distribution to compute the respective uncertainty. By applying the same uncertainty estimation method for both human and model assessments, a direct comparison of the results becomes possible.
@@ 3.2 Methods
Table 1: Task complexity based on human and model predictions. For a given input of class A, the prediction is correct if it is class A, or incorrect if it is any other class !A.
@@
A A A Easy
A A !A Medium
A !A A Difficult
A !A !A Difficult
@@ A A !A Medium
At last, to explore the potential of human insights to improve NNs uncertainty estimates, a subset of human annotations was incorporated into the model using soft labels while computing the CE loss. The quality of the updated model’s uncertainty estimates was then re-evaluated against the original human uncertainty annotations.
@@ 4 Results
This section reports the results obtained in this correlation study. Table 2 shows the correlation between human perceptual uncertainty and model uncertainty evaluated using the different techniques considered (namely, CE with MC dropout, EUAT, CALS, deep ensemble, isotonic regression, and DEUP) training a ResNet18 on Cifar10-H and CifarN, and a ResNet50 on ImageNet-16H.

### v2

@@ Abstract
Modern neural networks (NNs) often achieve high predictive accuracy but are poorly calibrated, producing overconfident predictions even when wrong. This miscalibration poses serious challenges in applications where reliable uncertainty estimates are critical. In this work, we investigate how human perceptual uncertainty compares to uncertainty estimated by NNs. Using three vision benchmarks annotated with both human disagreement and crowdsourced confidence, we assess the correlation between model-predicted uncertainty and human-perceived uncertainty. Our results show that current methods only weakly align with human intuition, with correlations varying significantly across tasks and uncertainty metrics. Notably, we find that incorporating human-derived soft labels into the training process can improve calibration without compromising accuracy. These findings reveal a persistent gap between model and human uncertainty and highlight the potential of leveraging human insights to guide the development of more trustworthy AI systems.
@@ 1 Introduction
Further, real-world applications, such as medical diagnostics, autonomous driving, and fraud detection, increasingly leverage model uncertainty to enhance trustworthiness by incorporating human oversight. In these systems, models can flag predictions with high uncertainty as requiring human review. This approach allows systems to balance automation and human judgment, ensuring that critical decisions are not solely reliant on potentially unreliable predictions. By raising alarms for ambiguous cases, these systems create a feedback loop where humans can validate, correct, or refine predictions, thus improving the model’s performance and fostering user trust. However, the deployment of uncertainty-based alarm systems creates a trade-off between the frequency of alarms raised and the associated costs. These costs can be reduced by aligning the model uncertainty estimation with human perception.
@@
To investigate this, we study three visual classification tasks that include human annotations of perceptual uncertainty, either through soft labels (reflecting inter-annotator disagreement) or explicit confidence scores. We compare these human-derived signals with uncertainty estimates from multiple NNs using metrics such as predictive entropy (PE) [36, 18]. We further test whether injecting human soft labels into training can improve the model’s alignment with human intuition and its calibration.
@@ 2 Related Work
A significant body of work has explored methods to enhance calibration of NNs. Early methods leveraged Bayesian Neural Networks (BNNs) [26, 21, 3, 39], which place a prior distribution over the model weights and infer a posterior given the data [33]. While theoretically sound, BNNs are computationally intensive and scale poorly to large models and datasets. To mitigate this, several approximate Bayesian methods have been proposed such as Monte Carlo (MC) Dropout [4] and Variational Inference (VI) [12]. Among these, MC Dropout, due to its higher efficiency, has become the most widely used technique for uncertainty estimation in NNs [22, 1, 4]. This technique estimates uncertainty by applying multiple dropout masks and performing multiple stochastic forward passes, effectively simulating an ensemble of subnetworks. The resulting predictions are aggregated using statistical measures, such as PE or variance, to quantify the model’s uncertainty.
In addition to uncertainty quantification, calibration has become a key property of NNs, as they are often overconfident when their predictions are incorrect [8]. To address this, a range of post-hoc calibration methods have been developed, including Platt Scaling [29], Isotonic Regression [41], Temperature Scaling [8], and Beta Calibration [15], which adjust model confidence scores after training. These techniques are computationally efficient and easy to apply, but do not alter the underlying uncertainty estimates. Consequently, their effectiveness can vary significantly across architectures and datasets [20], and they tend to degrade under distributional shifts, reducing reliability in out-of-distribution (OOD) scenarios [27, 8].
@@ 3 Human and Model Uncertainty Evaluation
Therefore, this work conducts a correlational study to compare human perception of uncertainty against uncertainty estimation outputted by NNs. The objectives are twofold: i) to evaluate how well model uncertainty aligns with human perceptual uncertainty, especially in ambiguous or uncertain scenarios, and ii) to explore whether model uncertainty can be enhanced or complemented by human decision-making. By identifying scenarios where human and model uncertainty assessments converge, diverge, or reveal poor calibration, we can tailor the training to develop more trustworthy models.
This study hypothesizes that human perceptual uncertainty strongly correlates with the uncertainty estimates generated by NNs. Specifically, we posit that as human uncertainty increases, model uncertainty should also increase. Furthermore, we hypothesize that the strength of this correlation might be dependent on task complexity, with correlations weakening as task complexity increases. Additionally, we propose incorporating human insights to enhance the quality of model uncertainty evaluations, thereby advancing the development of more trustworthy AI systems.
@@ 3.1 Experimental Setup, Benchmarks, and Baselines
3.1.1 Datasets and Models.
This work exploits publicly available datasets (namely, Cifar10-H [28], CifarN [40], and ImageNet-16H [37]) in the image recognition domain (normally used to train NNs) that include human annotations from multiple reviewers. The datasets containing human annotations have been labeled by multiple reviewers, with the number of reviews per image varying based on the benchmark used  [^1] ImageNet-16H and Cifar-N dataset received approximately four to six judgments per image, while Cifar-10H dataset has fifty annotations per image.. Further, ImageNet-16H extends the ImageNet [32] dataset by reducing the sample size to 4800 and limiting the number of classes to 16, while introducing noise in the images.
@@
To train the models, independently of the considered solution, we use stochastic gradient descent to minimize the Cross Entropy (CE) loss function using a learning rate of 0.1, a momentum of 0.9, and a batch size of 64 for all the models. To guarantee reproducibility, ensure a fair comparison, and mitigate the randomness inherent in training NNs, each method was trained using ten different seeds [17]. Throughout this study, when a subset of the dataset is required, stratified sampling techniques are applied to ensure that all classes are adequately represented and that the sample remains representative of the original dataset. Further, the models are evaluated on the same test set used for human annotations to ensure a consistent comparison between human and model uncertainty estimates. All models and training procedures were implemented in Python3 via the Pytorch framework and trained using a single Nvidia RTX A4000 GPU.
3.1.2 Baselines.
Furthermore, we resort to several state-of-the-art methods [16, 20, 22, 17, 8] that have been proposed to estimate model uncertainty. Moreover, one important foundation of all these works lies in the computation of uncertainty given a prediction. One of the most widely used techniques is MC Dropout [7], which provides a Bayesian approximation for uncertainty by sampling multiple dropout masks and aggregating the predictions using PE. Like the human uncertainty estimation, MC Dropout combined with PE does not separate epistemic from aleatoric uncertainty, so both human and model uncertainty will reflect the total uncertainty of a prediction. In this study, we compare different baselines to compute the model uncertainty. More in detail, we evaluate both post-processing calibration methods (resorting to Isotonic Regression [41], or DEUP [16]) and uncertainty-aware training algorithms (namely, CE [42], EUAT [22], CALS [20], or deep ensemble [17]).
@@
3.1.3 Evaluation Metrics.
We resort to the PE of the outputted distribution to compute the respective uncertainty. By applying the same uncertainty estimation method for both human and model assessments, a direct comparison of the results becomes possible.
@@ 3.2 Methods
Table 1: Task complexity based on human and model predictions. For each input belonging to class AA, the prediction is considered correct if classified as AA, and incorrect if classified as any other class ($\lnot A$).
@@
AA AA AA Easy
AA AA $\lnot A$ Medium
AA $\lnot A$ A Difficult
AA $\lnot A$ $\lnot A$ Difficult
@@
At last, to explore the potential of human insights to improve NNs uncertainty estimates, a subset of human annotations was incorporated into the model’s training using soft labels while computing the CE loss. The quality of the updated model’s uncertainty estimates was then re-evaluated against the original human uncertainty annotations.
@@ 4 Results
This section reports the results obtained in this correlation study. Table 2 shows the correlation between human perceptual uncertainty and model uncertainty evaluated using the different techniques considered (namely, CE with MC dropout, EUAT, CALS, deep ensemble, isotonic regression, and DEUP) training a ResNet18 on Cifar10-H and CifarN, and a ResNet50 on ImageNet-16H. In these experiments, human reviews are not incorporated into the training process.
@@ 6 Conclusion
Acknowledgments
This work was supported by the Fundação para a Ciência e a Tecnologia (Portuguese Foundation for Science and Technology) through the Carnegie Mellon Portugal Program under grant SFRH/BD/151470/2021, and by projects UIDB/50021/2020, C645008882-00000055.PRR and C628696807-00454142 (Center for Responsible AI), 101189689 (ACHILLES). This work was developed within the scope of the project no. 62 - “Responsible A”, financed by European Funds, namely "Recovery and Resilience Plan - Component 5: Agendas Mobilizadoras para a Inovação Empresarial", included in the NextGenerationEU funding program.